OpenStudy (anonymous):

What are the solutions to the equation |x - 10| - 4 = 2x? Help me solve this please

OpenStudy (anonymous):

= to 2x? how am I supposed to solve this.

OpenStudy (anonymous):

@abb0t

OpenStudy (anonymous):

\[|x-3|=14\\ x-3=14\\x-3=-14\] solve the bottom two

OpenStudy (anonymous):

one step each ad \(3\) to both sides clear?

OpenStudy (anonymous):

yes now I got it thank you c:

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

^ i needed help with that one . I posted the wrong attachment -.-

OpenStudy (xapproachesinfinity):

pls try solving that using what @satellite73 has taught you

OpenStudy (anonymous):

I had an answer for that question long ago.

OpenStudy (xapproachesinfinity):

First understand what abs value means?

OpenStudy (anonymous):

It could either equal to -x or x

OpenStudy (anonymous):

I just went through that lesson but I do not understand why it equals to 2x and in the lesson it did not teach how to get the x out of there.

OpenStudy (xapproachesinfinity):

okay good now what is |x-4|?

OpenStudy (xapproachesinfinity):

if you understand abs value you just need to do the definition again and again

OpenStudy (xapproachesinfinity):

|x-10|-4=2x don't bother yourself with that 2x first look at your left hand side First step isolate the abs value so it will be alone in the left hand side to do this add 4 to both sides

OpenStudy (anonymous):

added to 2x?

OpenStudy (xapproachesinfinity):

\(|x-10|-4=2x\\|x-10|=2x+4\)

OpenStudy (anonymous):

6x

OpenStudy (anonymous):

???

OpenStudy (xapproachesinfinity):

yes both side! when you do equation you if you add something to one side you must add it to the other so things stay right! this balancing do you the old scaling you put 4 kg of meat in one side while you have 1kg unit in the other so you add 3 more kg units to scale things this pretty much it

OpenStudy (anonymous):

I know those things, I just need to know why I have "=2x"

OpenStudy (xapproachesinfinity):

no no 2x+4 you can't go to say it is 6x okay 2x is different quantity that just 4 you have and x in there it is saying 2 of something which we don't know

OpenStudy (xapproachesinfinity):

friend! i told you don't pay attention to 2x that much okay focus on abs value

OpenStudy (xapproachesinfinity):

so far we are at this |x-10|=2x+4 now if 2x+4 is just some number what would you do x-10=2x+4 or -(x-10)=2x+4 right?

OpenStudy (anonymous):

yes

OpenStudy (xapproachesinfinity):

you can see that i have to part again x-10=2x+4 or -(x-10)=2x+4 just like with |x|=a===> x=a or -x=a

OpenStudy (xapproachesinfinity):

Okay now you have two equation that you need solve

OpenStudy (xapproachesinfinity):

okay show me that you can this level of algebra!

OpenStudy (xapproachesinfinity):

once you done! tag me to see your work